C. 7. Which of the following best describes the Junction Rule, one of Kirchoffs Rules? a. Current can only run out of a juncIf someone can answer any/all of these that would be much appreciated!
Answer #1

Solution:

Q: 7

Answer: d

Explanation: At a junction, there will be both currents entering into and currents leaving the junction.

Kirchhoff's Junction Laws states that " the total amount of current flowing into the junction equals the total amount of current leaving the junction."

Q: 8

Answer: d (no direction)

Explanation: The force on a charged particle moving in a magnetic field is given by F = qvBsin\theta ;

where  \theta is the angle between the direction of velocity(v) of the charged particle and the direction of magnetic field(B).

Here, the charged particle is moving antiparallel to the direction of magnetic field. So, the angle \theta = 180o.

As sin180o = 0, F = 0.

Hence, the charged particle does not experience any force in any direction.

Q: 9

Answer: d

Explanation: Only a changing magnetic flux linked with a coil, produces an induced current in the loop.

The flux linked with the coil can be changed with a varying magnetic field or due to the relative motion between the magnetic field and the coil.

Q: 10

Answer: e

Explanation: Since magnetic monopoles do not exist, the field cannot diverge from a point or converge to a point.

As we know, the magnetic field lines are continuous, closed loops. That means there is no starting or ending point of field line.

If we imagine a closed surface enclosing a bar magnet, the magnetic number of field lines entering the surface is equal to the number of field lines leaving it. So, there will be no net outflow or inflow of the field.

So, divergence of magnetic field does not happen.

Q: 11

Answer: c

Explanation: Electromagnetic wave is transverse in nature. Electric field, magnetic field, and the direction of propagation of wave, all three are mutually perpendicular. That means both electric and magnetic fields are perpendicular to each other, and the wave propagates in a direction perpendicular to both electric and magnetic fields.

For example, if electric field is along X direction and magnetic field is along Y direction, then the wave propagates(direction of velocity of wave) in the Z direction.

Q: 12

Answer: d

If the speaker increases the frequency, the wavelength decreases.

Explanation: The relation between wavelength(\lambda), frequency(\nu), and speed(v) of sound wave(or any wave), is

v = \nu \lambda

As the conditions in the air remain the same, the speed of sound wave remains constant.

Hence, to keep the wave speed constant, with the increase in frequency, the wavelength of the wave is decreased.